# Beyond Numeric Awards: In-Context Dueling Bandits with LLM Agents

### v1

@@ B.1.1 Design of Prompts
Figure 6: Original prompt for LLM-Env interaction in dueling bandit setting with temperature = 0, including context $C$, summarized history $H_{t}$, and zero-shot chain-of-thought (CoT) reasoning instructions $R$ (see Section 2).
@@
Figure 7: Perturbed prompts used to evaluate the robustness of IF-Enhanced LLM. The Biased History prompt (Left) injects an incorrect dueling history favoring non-optimal arms, while the Reversed Goal prompt (Right) adversarially modifies the objective from maximizing to minimizing reward. Both prompts maintain the zero-shot chain-of-thought (CoT) reasoning, and temperature setting of 0.
B.1.2 Exemplars of GPT-4 Turbo
We provide an example of five consecutive time steps using the original prompt (see Figure 6) to demonstrate GPT-4 Turbo’s decision-making process in the dueling bandits setting. This example highlights GPT-4 Turbo’s capability to analyze available information, reason about the relative strengths of the arms, and make informed decisions on which pair of arms to duel next.
Answer at Time Step 0: 1. Analysis: Since no duels have been conducted yet, all empirical probabilities between arms are at their initial value of 0.50, indicating no preference or historical data to suggest one arm is better than another. To begin identifying the Condorcet Winner, we need to start collecting data by conducting duels between different arms. A balanced approach would be to start dueling pairs that have not yet competed, but since all pairs are at the same starting point, we can choose any pair. 2. Next Pair: (1, 2)
Answer at Time Step 1: 1. Analysis: From the results provided, the only conclusive duel so far is between Arm 1 and Arm 2, where Arm 1 won, giving P(1 > 2) = 1.00. This suggests that Arm 1 might be stronger than Arm 2, but more data is needed to confirm its strength relative to other arms. All other pairwise comparisons have not yet resulted in any duels, so their probabilities remain at the initial estimate of 0.50, indicating no preference between the arms based on the data. To better understand the relative strengths of the arms and potentially identify the Condorcet Winner, it is crucial to increase the number of duels among arms with less information. 2. Next Pair: Given that Arm 1 has shown some potential by winning against Arm 2, it would be informative to test Arm 1 against another arm. Since Arm 1 has not yet been tested against Arm 3, Arm 4, or Arm 5 in a conclusive duel, any of these could be a good choice. However, to balance the exploration among other arms and gather more data on their performance, let’s choose a pair that has not been tested yet. Arm 3 and Arm 4 have not dueled conclusively, and understanding their relative strength could provide insights into the overall hierarchy of the arms. Choose: (Arm 3, Arm 4)
Answer at Time Step 2: 1. Analysis: - Arm 1 vs Arm 2: Arm 1 has a clear advantage over Arm 2 with a win probability of 1.00. - Arm 3 vs Arm 4: Arm 3 has a clear advantage over Arm 4 with a win probability of 1.00. - Other Duels: All other duels have resulted in a 50% win probability for each arm, indicating no clear advantage based on the limited data available. Given these results, Arm 1 and Arm 3 appear to be stronger contenders in their respective duels. However, more data is needed to confidently identify the Condorcet Winner. To gather more information, it would be strategic to pit arms against each other where the win probabilities are still uncertain (i.e., 0.50), especially focusing on those arms that have shown potential strength. 2. Next Pair: (Arm 1, Arm 3)
Answer at Time Step 3: 1. Analysis: - Arm 1 has shown strong performance against Arm 2 and Arm 3, winning all duels against them. However, it has not yet won against Arm 4 and Arm 5, with the results still uncertain (P=0.50). - Arm 3 has won against Arm 4, suggesting it might be stronger than Arm 4. - Arms 4 and 5 have not won any duels yet, and their probabilities against other arms are still at 0.50, indicating uncertainty and lack of data. - Arms 2, 4, and 5 have not won any duels, and their probabilities against other arms are mostly at 0.50, indicating a need for more data to determine their strengths. Given this analysis, it seems prudent to focus on gathering more information about the arms that have uncertain outcomes and have not shown strong winning probabilities. Specifically, more data is needed for Arm 1 against Arms 4 and 5, as well as between Arms 4 and 5 themselves, to better understand their relative strengths. 2. Next Pair: (Arm 1, Arm 4)
Answer at Time Step 4: 1. Analysis: - Arm 1 has shown strong performance against Arms 2, 3, and 4, winning all duels against them. However, it has not yet won against Arm 5, with the probability of winning currently estimated at 0.50. - Arm 3 has also shown some strength by winning against Arm 4. - The other arms (2, 4, 5) have not won any duels yet, and their probabilities of winning against each other are all estimated at 0.50, indicating no clear evidence of superiority among them. - Given the current data, Arm 1 appears to be a strong candidate for the Condorcet Winner, but more evidence is needed, particularly in its performance against Arm 5. To further clarify the potential of Arm 1 and to explore the capabilities of Arm 5, which has not lost yet but also has not won (except in an undecided duel against Arm 1), it would be strategic to have them duel again. This will help in determining if Arm 1 can consistently outperform Arm 5 or if Arm 5 has potential that hasn’t been fully revealed yet. 2. Next Pair: (Arm 1, Arm 5)
B.2 Supplementary Experiments
B.2.1 Comparisons with Different Metrics
@@
We present supplementary results to complement our case studies. (i) Figure 8 shows the strong and weak regret comparisons for the Hard instance. (ii) Figure 9 introduces the Best Arm Ratio, measuring how often the algorithms select the best arm (Condorcet winner) in a duel for both instances. (iii) Figure 10 examines the generalized variance of the strong and weak regret for both instances.
@@
Figure 8: Comparisons between LLM agents and various classic DB algorithms. Left and Right: strong and weak regret for the Hard instance. Results for the Easy instance is presented in Figure 2.
@@
Figure 9: The Best Arm Ratio represents the fraction of duels that include the best arm (Condorcet winner) over different time intervals in the Easy instance (Left) and the Hard instance (Right). This metric measures how often an algorithm selects the best arm to participate in a duel. A higher Best Arm Ratio indicates that the algorithm is more effective at identifying and exploiting the best arm. Three LLMs (GPT-3.5 Turbo, GPT-4, GPT-4 Turbo) and two state-of-the-art baselines (Self-Sparring and DTS) are compared against each other.
@@
Figure 10: Comparison of the generalized variance of strong and weak regret between LLMs and baseline algorithms on the Easy (Left) and Hard (Right) instances. In the Easy instance, GPT-4 Turbo exhibits the lowest average generalized variance. For the Hard instance, GPT-4 Turbo maintains a variance level comparable to state-of-the-art baseline algorithms (except BTM and SAVAGE, which are in an early exploration stage).
B.2.2 Duel Selection Trajectory
@@
LLM Agents and Baseline Algorithms
Figure 11: Comparison of duel selection trajectories among GPT-4 Turbo, Self-Sparring, and DTS on the Easy (Top Row) and Hard (Bottom Row) instances. The decision trajectories of GPT-4 Turbo exhibit a clear pattern of continuous exploration without converging to the best arm. In contrast, Self-Sparring and DTS demonstrate more structured exploration patterns and convergence properties on both instances.
Figure 12: Local optima trajectories of GPT-3.5 Turbo (Left), GPT-4 (Middle), and GPT-4 Turbo (Right, with noisy prompt) on the Hard instance. Less capable LLMs, such as GPT-3.5 Turbo and GPT-4, could get stuck comparing suboptimal arms on hard preference structure. Even for GPT-4 Turbo, noisy prompts with biased history (see Figure 7) can lead it to be trapped in bad tournaments.
Figure 13: Comparison of success (Left) and failure (Right) cases for the Convergence-Triggered GPT-4 Turbo intervention strategy discussed in Section 5.1. While it works for most cases due to GPT-4 Turbo’s strong capability (Left), sometimes this naive intervention can reinforce suboptimal choices (Right) on the Hard instance.
IF-Enhanced LLM
Figure 14: Trajectory of the classic IF2 algorithm with different initial incumbent arm selections. The left plot shows the behavior of IF2 when a good initial incumbent arm is randomly selected, leading to faster convergence and better overall performance. In contrast, the right plot illustrates the case where a bad initial incumbent arm is chosen, resulting in slower convergence and suboptimal performance.
Figure 15: Duel selection trajectory comparison between GPT-4 Turbo and IF-Enhanced LLM under different prompt settings (see Figures 6 and 7). (i) Top Row: With the original prompt, IF-Enhanced LLM leverages GPT-4 Turbo’s exploration ability and guarantees convergence through the IF2 phase. (ii) Middle Row: With a noisy prompt (biased history), IF-Enhanced LLM overcomes the limitation of standalone GPT-4 Turbo getting trapped in local optima by employing uniform comparisons in the IF2 phase. (iii) Bottom Row: With an adversarial prompt (reversed goal), IF-Enhanced LLM maintains robust behavior despite the adversarial modification. Across all scenarios, IF-Enhanced LLM demonstrates superior performance and robustness compared to standalone GPT-4 Turbo.

### v3

@@ C.1.2 Design of Prompts
Figure 7: Original prompt for LLM-Env interaction in dueling bandit setting with temperature = 0 (except o1-preview, which is in beta phase, its system prompt and user prompt are concatenated together with a fixed temperature = 1), including context $P$, summarized history $H_{t}$, and zero-shot chain-of-thought (CoT) reasoning instructions $R$ (see Section 2).
@@
Figure 8: Perturbed prompts used to evaluate the robustness of LEAD. The Biased History prompt (Left) injects an incorrect dueling history favoring non-optimal arms, while the Reversed Goal prompt (Right) adversarially modifies the objective from maximizing to minimizing reward. Both prompts maintain the zero-shot chain-of-thought (CoT) reasoning and temperature setting as before.
C.1.3 Exemplars of GPT-4 Turbo and o1-preview
We present exemplars using the original prompt (see Figure 7) to illustrate the decision-making process of both GPT-4 Turbo and o1-preview in the dueling bandits setting. These examples highlight how each model interprets the available data, reasons about the relative strengths of the arms, and selects the next pair of arms to duel. The exemplars show both successful convergence cases and the failures of each model.
Systematic Biases
Biased Understanding of GPT-4 Turbo and o1-preview System Prompt: You are a Dueling Bandit algorithm with 5 arms. Each pair of arms (i, j) has an unknown probability $P(i>j)$ of arm $i$ winning against arm $j$ in a pairwise duel. There exists a best arm (Condorcet Winner) among the 5 arms, which has a greater than 50% probability of winning against any other arm in a pairwise duel. At each time step, I will provide you with a summary that shows the past dueling results and empirical probabilities. Then you must choose 2 arms to duel against each other. The reward you receive is the sum of the unknown probabilities of the two chosen arms beating the best arm. The maximum reward in each time step is achieved if the best arm duels against itself. Your goal is to maximize the cumulative sum of rewards of the chosen two arms over a given time horizon $T=1000$. Question: Can an arm duel against itself? Answer Exemplars: Example 1: In a traditional dueling bandit problem, an arm dueling with itself doesn’t make practical sense because the outcome is trivially known—it would always win. The typical setup for dueling bandits is to choose two distinct arms to engage in a duel, which provides meaningful information about their relative strengths. Example 2: In the context of a dueling bandit problem, an arm dueling with itself typically doesn’t make sense because the purpose of a duel is to compare two different arms. Dueling an arm with itself wouldn’t provide any new information about the relative strengths of the arms, as the outcome is trivially known: the arm would win against itself 100% of the time. Example 3: In the context of a dueling bandit problem, an arm cannot duel with itself because the concept of a duel inherently involves comparing two different entities. The purpose of a duel is to establish which of the two arms is superior, or to collect comparative information. An arm dueling itself would not generate any valuable information or result, as it would trivially win against itself. Therefore, you need to choose two distinct arms for each duel in a dueling bandit problem. Example 4: In the context of a dueling bandit problem, an arm cannot duel with itself because the concept of a duel inherently involves comparing two different entities. The purpose of a duel is to establish which of the two arms is superior, or to collect comparative information. An arm dueling itself would not generate any valuable information or result, as it would trivially win against itself. Therefore, you need to choose two distinct arms for each duel in a dueling bandit problem. Comment: LLMs lack a fundamental understanding of the DB problem
Success cases
Successful Convergence Case of GPT-4 Turbo 1. Analysis: - Arm 5 appears to be the strongest contender, showing a high probability of winning against all other arms except Arm 2, where it still maintains a majority win rate. - Arm 3 has a strong win rate against Arm 4 and a reasonable rate against Arm 2, but significantly loses to Arm 5. - Arm 2 has a mixed performance, winning against Arm 1 and Arm 4 but losing to Arm 5 and Arm 3. - Arm 1 and Arm 4 are the weakest, with Arm 1 losing to almost all arms and Arm 4 only winning against Arm 1. - Given the goal to maximize the cumulative sum of rewards, focusing on duels involving Arm 5 seems most beneficial, as it has the highest probability of winning against other arms, suggesting it might be the Condorcet Winner. 2. Next Pair: (Arm 5, Arm 5). - Dueling Arm 5 against itself maximizes the reward for this step, as it likely has the highest unknown probability of beating the best arm (itself in this case). This strategy ensures the highest possible reward in a single duel under the given conditions. Comment: Successfully understand the problem and suggest a converged arm pair
Successful Convergence Case of o1-preview 1. Analysis: Given the dueling results so far, we are tasked with maximizing the cumulative reward over the remaining time horizon $T=2000$ steps. The reward at each time step is the sum of the unknown probabilities of the two chosen arms beating the best arm (the Condorcet winner). To maximize our reward, we need to identify the arm that is most likely the best based on the empirical data and choose it to duel against itself. This is because the maximum reward at each time step is achieved when the best arm duels against itself. Analyzing the empirical probabilities: - Arm 5 has beaten all other arms with a probability greater than 50%: - $P(5>1)=0.75$ $P(5>2)=0.75$ $P(5>3)=0.64$ $P(5>4)=1.00$ This suggests that Arm 5 is the most likely Condorcet winner based on the data. Therefore, to maximize the cumulative reward, we should choose Arm 5 to duel against itself, as this maximizes the sum of probabilities $2\times P(5>\text{BestArm})$. 2. Next Pair: (Arm 5, Arm 5) Comment: Successfully understand the problem and suggest a converged arm pair
Failure cases
Convergence Instability Case of GPT-4 Turbo: 1. Analysis: - Arm 5 appears to be a strong contender as it has a high winning probability against most other arms, particularly against Arm 1, Arm 2, Arm 3, and Arm 4. Notably, it has a significant number of duels against Arm 3, which provides a robust estimate of its strength relative to Arm 3. - Arm 3 also shows strength, especially against Arm 4 and Arm 2, but it has a lower winning probability against Arm 5. - Arm 1 and Arm 4 seem to be weaker overall, with Arm 1 losing all duels against Arm 5 and Arm 4 having a low win rate against Arm 5. - Arm 2 is somewhat in the middle, with a mixed record against the other arms. Given the results, Arm 5 is likely the Condorcet Winner as it has a greater than 50% win rate against all other arms. To maximize the reward, it would be beneficial to continue testing Arm 5 against other strong contenders to confirm its dominance and further understand the dynamics between the other arms. 2. Next Pair: - This pairing is chosen because while Arm 5 has shown strong performance against Arm 3, the large number of duels suggests that further testing could provide additional insights and confirm stability in the empirical probabilities. Additionally, dueling these two strong arms could potentially yield a high reward if Arm 5 continues to dominate. Choose: (Arm 5, Arm 3) Comment: Keeps comparing the same top arms without committing to a single winner
Biased Understanding Case of o1-preview: 1. Analysis: From the given data, we can infer that Arm 5 is likely the best arm (Condorcet Winner) because it has beaten all other arms with probabilities greater than 50 - P(5 > 1) = 0.75 - P(5 > 2) = 0.75 - P(5 > 3) = 0.66 - P(5 > 4) = 1.00 Our goal is to choose two arms (armA and armB) such that the sum P(armA > best arm) + P(armB > best arm) is maximized. From the empirical probabilities: - P(3 > 5) = 0.34 - P(2 > 5) = 0.25 - P(1 > 5) = 0.25 - P(4 > 5) = 0.00 Arm 3 has the highest probability of beating Arm 5, followed by Arms 2 and 1. Arm 4 has a zero probability and should not be selected. Between Arms 1 and 2, Arm 2 has participated in more duels against Arm 5, providing a more reliable estimate of its probability to beat the best arm. Additionally, Arm 2’s performance against other arms suggests it may be slightly better than Arm 1. Therefore, the best choice to maximize our expected reward is to select Arm 3 and Arm 2. 2. Next Pair: (Arm 3, Arm 2) Comment: Exhibits a biased understanding of the problem description: while successfully identifying the Condorcet Winner, it still selects suboptimal arms in duel.
C.2 Supplementary Experiments
C.2.1 Comparisons with Different Metrics
@@
We present supplementary results to complement our case studies. (i) Figure 9 shows the strong and weak regret comparisons for the Hard instance. (ii) Figure 12 presents comparisons under different numbers of arms $K$, illustrating the impact of the Relative Decision Window. (iii) Figure 13 introduces the Best Arm Inclusion Ratio and the Converged Best Arm Ratio. (iv) Figure 14 examines the generalized variance of the strong and weak regret for both instances.
@@
Figure 9: Comparisons between LLM agents and various classic DB algorithms. Left and Right: strong and weak regret for the Transitive-Hard instance. Results for the Transitive-Easy instance is presented in Figure 2. We evaluate only three LLMs on the Transitive-Hard instance due to our research goals and high API costs: (i) The results for the Transitive-Hard instance are qualitatively similar to those for the Transitive-Easy instance; (ii) Obviously, the Transitive-Easy instance offers higher distinguishability, allowing us to observe convergence and regret differences within a feasible number of steps.
Figure 10: Comparisons between GPT-4 Turbo and various classic DB algorithms. Left and Right: strong and weak regret for the Intransitive-Easy instance. Results for the Intransitive-Hard instance is presented in Figure 11. We evaluate only our top-performing LLM on the Intransitive-Easy and Intransitive-Hard instance to examine the scalability limitation.
Figure 11: Comparisons between GPT-4 Turbo and various classic DB algorithms. Left and Right: strong and weak regret for the Intransitive-Hard instance.
Figure 12: Cumulative strong and weak regret comparisons between LLM agents and classic dueling bandit algorithms on Transitive-Easy instance under different numbers of arms K. Top Left and Top Right: K=5, where GPT-4-Turbo significantly outperforms other methods on weak regret. Bottom Left and Bottom Right: K=10, where the performance of GPT-4-Turbo degrades as the number of arms increases.
@@
Figure 13: Four LLMs (GPT-3.5 Turbo, GPT-4, GPT-4 Turbo, o1-preview) and two state-of-the-art baselines (Self-Sparring and DTS) are compared against each other on the Transitive-Easy instance over different time intervals. Left: the Best Arm Inclusion Ratio represents the fraction of duels that include the best arm (Condorcet winner). Middle: the Converged Best Arm Ratio represents the proportion of duels where the best arm duels against itself for exploitation. Right: the Suboptimal Duel Ratio represents the proportion of duels where both arms selected in duel are suboptimal arms. We observed that while o1-preview can transit from exploration to exploitation (high Converged Best Arm Ratio), it selects more optimal arms (high Suboptimal Duel Ratio) due to the reinforced biased understanding as discussed in Section 3.2.
@@
Figure 14: Comparison of the generalized variance of strong and weak regret between three LLMs and baseline algorithms on the Transitive-Easy (Left) and Transitive-Hard (Right) instances. In the Easy instance, GPT-4 Turbo exhibits the lowest average generalized variance. For the Transitive-Hard instance, GPT-4 Turbo maintains a variance level comparable to state-of-the-art baseline algorithms (except BTM and SAVAGE, which are in an early exploration stage).
C.2.2 Duel Selection Trajectory
@@
Figure 15: Comparison of duel selection trajectories between GPT-4 Turbo (Left) and o1-preview (Right) on the Transitive-Easy instance. GPT-4 Turbo achieves low weak regret by consistently selecting the best arm, though it struggles to converge to a single best arm. In contrast, o1-preview shows better convergence behavior, but its weak regret performance is worse than GPT-4 Turbo due to incomplete or biased understanding, as illustrated by the o1-preview exemplar in Appendix C.1.3.
Figure 16: Local optima trajectories of GPT-3.5 Turbo (Left), GPT-4 (Middle), and GPT-4 Turbo (Right, with noisy prompt) on the Transitive-Hard instance. Less capable LLMs, such as GPT-3.5 Turbo and GPT-4, could get stuck comparing suboptimal arms on hard preference structure. Even for GPT-4 Turbo, noisy prompts with biased history (see Figure 8) can lead it to be trapped in bad tournaments.
Figure 17: Comparison of success (Left) and failure (Right) cases for the Convergence-Triggered GPT-4 Turbo intervention strategy discussed in Section 4.1. While it works for most cases due to GPT-4 Turbo’s strong capability (Left), sometimes this naive intervention can reinforce suboptimal choices (Right) on the Transitive-Hard instance.